# wd1: Weighted Policy Optimization for Reasoning in Diffusion Language Models

### v1

@@ B.2 Reward Function
We use the same reward function as in d1 (Zhao et al., 2025). For completion, we provide the details as following.
GSM8K. Following the Unsloth reward setup [^8] https://unsloth.ai/blog/r1-reasoning, we apply five addtive components: XML Structure Reward: +0.125 per correct tag; small penalties for extra content post-tags. Soft Format Reward: +0.5 for matching the pattern
@@
Strict Format Reward: +0.5 for exact formatting with correct line breaks. Integer Answer Reward: +0.5 if the answer is a valid integer. Correctness Reward: +2.0 if the answer matches ground truth.
@@ B.3 Sampling from Geometric Mixture
Although the approximate sampling strategy introduced in Section 3.2 eliminates the need to approximate the reference policy’s likelihood, it incurs computational overhead, as generating a full completion requires multiple forward passes through the dLLM—compared to a single pass for likelihood estimation. An alternative is to sample from $\pi_{\text{old}}$ and shift the advantage to $\hat{A}_{i}=A^{\pi_{\text{old}}}(q,o_{i})+\beta\log\hat{\pi_{\text{ref}}}/(\lambda+\beta)$, which reintroduces the need for reference policy likelihood approximation. However, our method disables reference policy regularization, following prior works, thereby avoiding this issue entirely.
@@ B.5 Training Cost Estimation
For the runtime measurements reported in Table 2, we set $\mu=8$ and train for a total of 6 global steps, corresponding to 48 gradient update steps. We use a batch size of 4 and the rest of the hyperparameters are the same as in Table 5. To estimate the number of function evaluations (NFEs) involved in computing likelihood approximations, we count only the forward passes, as the number of backward passes remains consistent across methods. The additional NFEs observed in the d1 model arise from evaluating the likelihood under both the old and reference models, which are used for regularization. These extra evaluations are required only when new samples are drawn, as their outputs can be cached and reused across all gradient updates for $\mu$. We additionally report the number of floating-point operations (FLOPs) per global training step, measured using the Flops Profiler from Rasley et al. (2020).
@@ Appendix C Additional Experiments
We additionally report results for comparison to the results of the baseline d1 reported in the paper (Zhao et al., 2025). As shown in  Table 6, our method wd1 evaluated and selected from less checkpoints, can outperform d1 with a large margin in Sudoku and Countdown, achieving comparable performance in math problem-solving tasks.
@@
+ diffu-GRPO (reproduced) 16.1 9.3 27.0 34.0 80.7 79.1 34.4 39.0
@@
wd1 25.2 24.2 51.2 46.1 80.8 82.3 34.4 39.0
wd1-SFT 26.5 24.2 43.4 43.4 80.7 82.0 36.4 39.0
wd1-P 6.69 6.84 13.67 4.69 65.66 78.17 29.40 22.80
Table 6: Test accuracy ($\uparrow$) across different tasks. Our method demonstrates higher accuracy, especially significant in Sudoku and Countdown. The shaded numbers indicate where our method performs the best.
We additionally provide reward dynamics in comparison to wd1-SFT in training. In Sudoku and Countdown, directly training with wd1 without SFT demonstrate significantly more efficient and stable learning process. In GSM8k and MATH500, the different is negligible.

### v2

@@ B.2 Reward Function
To train wd1 and reproduce d1, we use the reward function defined in (Zhao et al., 2025). For completion, we provide the details as following.
@@
To train wd1++, we leverage Math-Verify (Kydlíček, ), constructing a simple verifier reward function to evalaute on GSM8K and MATH500.
@@ B.3 Sampling from Geometric Mixture
Although the sampling strategy eliminates the need to approximate the reference policy’s likelihood, it incurs computational overhead, as generating a full completion requires multiple forward passes through the dLLM—compared to a single pass for likelihood estimation. An alternative is to sample from $\pi_{\text{old}}$ and shift the advantage to $\hat{A}_{i}=A^{\pi_{\text{old}}}(q,o_{i})+\beta\log\hat{\pi_{\text{ref}}}/(\lambda+\beta)$, which reintroduces the need for reference policy likelihood approximation. However, policy ratio has been removed, and the reference model can be reused when conducting multiple gradient updates with the same batch of rollouts (off-policy). The increased computational burden is slight.
@@ B.5 Training Cost Estimation
For the runtime measurements reported in Table˜2, we set $\mu=8$ and train for a total of 6 global steps, corresponding to 48 gradient update steps. We use a batch size of 4 and the rest of the hyperparameters are the same as in Table˜6. To estimate the number of function evaluations (NFEs) involved in computing likelihood approximations, we count only the forward passes, as the number of backward passes remains consistent across methods. The additional NFEs observed in the d1 model arise from evaluating the likelihood under both the old and reference models, which are used for regularization. These extra evaluations are required only when new samples are drawn, as their outputs can be cached and reused across all gradient updates for $\mu$. We additionally report the number of floating-point operations (FLOPs) per global training step, measured using the Flops Profiler from Rasley et al. (2020).
B.6 Computing Resources
For both wd1 and d1, RL training is conducted on four NVIDIA A100 GPUs (80GB), and SFT is performed on four A6000 GPUs (48GB). For wd1++ and MDPO, RL training is conducted on $\timesA8$00 (80GB).
@@ Appendix C Additional Experiments
We additionally report results for comparison to the results of the baseline d1 reported in the paper (Zhao et al., 2025). As shown in  Table˜7, our method wd1 evaluated and selected from less checkpoints, can outperform d1 with a large margin in Sudoku and Countdown, achieving comparable performance in math problem-solving tasks.
C.1 Summary of wd1  Results
Table 7: Test accuracy across different tasks. Our method demonstrates higher accuracy, especially significant in Sudoku and Countdown. The shaded area indicates where our method outperforms.
@@ C.1 Summary of wd1  Results
+ diffu-GRPO (reproduced) 16.1 11.7 27.0 34.0 80.7 79.1 34.4 39.0
@@
wd1 76.4 62.8 51.2 46.1 80.8 82.3 34.4 39.0
Table 8: Training cost. The training steps to obtained the best post-trained model of three methods are 20, 150, and 7500. To compute the total rollouts, we need to compute the average rollouts in a single training step. Gradient steps per rollout batch represents the number of gradient descent conducted with a single batch of rollouts. In other words, 1 represents it is a pure on-policy RL training, and for any value >12>12, off-policy RL is executed. Total Batch Size is computed by multiplying per-device batch size, gradient accumulation and the number of gpus. Therefore, the average number of rollouts used for single step gradient descent should be computed by total batch size divided by gradient steps per rollout batch.
Hyperparameter wd1++ MDPO d1
Training step of the best checkpoint 20 150 7500
Training Steps per Rollout batch 1 1 12
Per-Device Batch Size 4 1 6
Gradient Accumulation 2 16 2
GPUs used for training 8 8 4
Total Batch Size 64 128 48
Avg. Rollouts per Step 64 128 4
Total Rollouts 1280 19200 30000
We additionally provide reward dynamics in comparison to wd1-SFT in training. In Sudoku and Countdown, directly training with wd1 without SFT shows significantly more efficient and stable learning process. In GSM8k and MATH500, the difference is negligible.
@@
C.2 Additional Ablation Study
Table 9: Ablation on weight $\lambda$ to combine positive $w^{+}$ and negative weights $w^{-}$ in wd1 on Sudoku. Specifically, the final weight assigned to log-likelihood is computed as $\lambda w^{+}+(1-\lambda)w^{-}$.
Combined Weight Accuracy Effective Tokens
0.5 25.63% 326.97
0.4 11.77% 240.04
0.6 14.11% 220.13
We provide additional ablation study on the combined weight to confirm our analysis that the positive and negative samples terms in the loss function should be assigned equal proportion, due to the side case of a batch of all-negative generated responses (see the paragraph below Equation˜9). Assigning equal proportions to positive and negative weights is not arbitrary but rather the most robust design. This can be understood through two critical failure modes that arise from imbalanced proportions:
  • When positive weight has larger proportion: In scenarios where all sampled completions have uniformly low rewards, a larger proportion of positive weights would paradoxically increase the log-likelihood of negative samples during wd1 optimization, which is clearly undesirable and contradicts the learning objective.
  • When negative weight has larger proportion: Conversely, when all generated completions achieve uniformly high rewards, an insufficient proportion of positive weights would result in unlearning high-quality samples.
To empirically validate this analysis, we conducted experiments on the Sudoku dataset with varying mixing proportions. The results, presented in the table below, confirm our theoretical predictions.
Figure 4: Left: Ablation study on $\psi$ in the weights (Equation˜9) on GSM8K. Right: wd1 training on MATH with a random seed different from the seed used in our main experiments. The abrupt decrease of the rewards in the early training (see Figure˜3 MATH500) disappears.
In all benchmark evaluations, we fix the hyperparameter $\psi=1$, which controls the scale of the exponential weighting in wd1. To validate this choice, we provide an ablation study on the coefficient $\psi=1$ in the exponential weight of wd1 (Equation 9) below. Larger values $\psi$ leads to more extreme weight assigned to the samples. According to Figure˜4, the training of applying different $\psi$ converges to similar rewards if $\psi$ is small. Overly large value (e.g. 10) can cause performance drop, implying that extreme weight assignment is detrimental.
C.3 Coding Benchmarks
We conduct 200 steps of wd1 training on AceCode-87K (Zeng et al., 2025) following the implementation of Open-R1 (Hugging Face, 2025). Our method achieves consistent improvements over the base model.
Table 10: Comparative performance of wd1 improvements compared to base model LLaDA-8B-Instruct. We present the results of wd1 fine-tuned with AceCode dataset (Zeng et al., 2025).
Task Gen Length Steps Block Size wd1 LLaDA
HumanEval 256 128 32 34.76 (+3.66) 31.10
HumanEval 256 256 32 39.02 (+1.82) 37.20
HumanEval 512 512 32 36.59 (+0.61) 35.98
MBPP 128 128 32 39.2 (+2.40) 36.8
MBPP 256 256 32 36.6 (+1.00) 35.6
MBPP 512 512 32 36.8 (+0.40) 36.4
C.4 Training Dynamics
Figure˜3 presents the reward dynamics over gradient steps during training. wd1 exhibits a notably faster reward increase compared to d1, highlighting its superior sample efficiency–effectively leveraging the reward signal to accelerate policy optimization. In addition, Figure˜5 shows the average length of generated completions during training. On math reasoning benchmarks such as GSM8K and MATH500, wd1 converges to shorter output sequences than d1, suggesting improved token efficiency while maintaining or improving performance.
Appendix D Limitations
Similar to other RL-based approaches, wd1 may lose effectiveness when all generations within a sampled group receive identical rewards. This situation can occur under several conditions—for example, when the training dataset is either too simple or too challenging for the base model. Nonetheless, such cases can be mitigated through careful reward design and the incorporation of curriculum learning strategies.
An additional limitation of this work is that the current wd1 framework is restricted to text-based reasoning. Extending it to multimodal reasoning or unified diffusion-based models (e.g., (Yang et al., 2025)) represents a valuable direction for future research.
A final limitation concerns the likelihood approximation used in wd1. Our approach relies on the d1-based approximation, which is computationally efficient but introduces bias. Although some prior works employ ELBO-based estimators (e.g., DCE), they require additional computational overhead (Zhao et al., 2025) often exhibit high variance, as demonstrated in Figure 1. This trade-off highlights an important area for further exploration.
D.1 Additional Analysis on Unlearning
We provide extended demonstrations for Remark Remark˜2, focusing specifically on the theoretical insights underlying the interpretation of the negative-sample reinforcement term in wd1 as a form of data unlearning. Under the DCE likelihood approximation, the negative-sample reinforcement term in wd1 becomes
$\mathbb{E}_{o\sim p^{\prime}_{0}(\cdot)}\bigg[w^{-}(q,o)\cdot\log\pi_{\theta}(o|q)\bigg]=\mathbb{E}_{o\sim p^{\prime}_{0}(\cdot)}\bigg[\exp(-A(x_{0}))\cdot\log\pi_{\theta}(o|q)\bigg]$ (67)
$=$ $\mathbb{E}_{x_{0}\sim p^{\prime}_{0}(\cdot)}\bigg[\exp\big(-A(x_{0})\big)\cdot\underbrace{\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x_{0})}\big[\sum_{x^{i}_{t}=[\texttt{mask}]}-\frac{1}{t}\log p_{\theta}(x_{0}^{i}|x^{\text{UM}}_{t})\big]}_{\mathcal{L}_{\text{DCE}}}\bigg]$ (68)
$=$ $\mathbb{E}_{x^{-}_{0}\sim p_{\text{data}}}\bigg[\underbrace{\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x^{-}_{0})}\big[\sum_{x^{i}_{t}=[\texttt{mask}]}-\frac{1}{t}\log p_{\theta}(x_{0}^{-,i}|x^{\text{UM}}_{t})\big]}_{\mathcal{L}_{\text{DCE}}\ \Leftrightarrow\ \text{ELBO}}\bigg],$ (69)
where $p_{\text{data}}(x^{-}_{0})=p^{\prime}_{0}(x^{-}_{0})\frac{\exp(-A(x^{-}_{0}))}{\sum_{x^{-}_{0}}p^{\prime}_{0}(x^{-}_{0})\exp(-A(x^{-}_{0}))}$. Equation˜69 holds by simply applying importance sampling.
Since DCE is equivalent to the evidence lower bound (ELBO) of masked discrete diffusion models, we draw an analogy between the final objective in Equation˜69 and data unlearning in diffusion models (Alberti et al., 2025). Equation˜69 can be viewed as a direct masked discrete–diffusion extension of NegGrad (Golatkar et al., 2020), which aims to minimize the evidence lower bound of the log-likelihood on samples with lower advantage.